OpenStudy (anonymous):

A batter hits a baseball upward with an initial speed of 96 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt − 16t2 where h represents height in feet and r represents the rate in feet per second.

OpenStudy (anonymous):

in this example \(r=96\) although usually it is called \(v_0\)

OpenStudy (anonymous):

i guess the batter hits the ball that is lying on the ground, because there is supposed to be an initial height involved as well no matter your job is to solve \[96t-16t^2=0\] so you know how to solve it?

OpenStudy (anonymous):

"no" is an ok answer, if you do not now how i can walk you through it

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok fine

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[96t-16t^2=0\] is what we have to solve it is probably easier to multiply both sides by \(-1\) i.e. change each sign and solve \[16t^2-96t=0\]

OpenStudy (anonymous):

\(16t^2\) and \(96t\) both have a common factor of \(t\) they also both have a common factor of \(16\) because \(16\times 6=96\)

OpenStudy (anonymous):

so you can factor this as \[16t(t-6)=0\]

OpenStudy (anonymous):

so t=6?

OpenStudy (anonymous):

one answer is \(16t=0\iff t=0\) but that is not your answer, that just says at the beginning the ball is on the ground the other is right

OpenStudy (anonymous):

\(t=6\) is the answer

OpenStudy (anonymous):

ty!

OpenStudy (anonymous):

yw